Q1. Express the following equation in slope-intercept form: 13x - 3y = 18
Q2.Express the following equation in slope-intercept form: 17x - 10y = 80
Q3.Express the following equation in slope-intercept form: 5x + 8y = 40

Answer by BigToosie(32)

Q1. Express the following equation in slope-intercept form: 13x - 3y = 18
slope-Intercept form y = mx + b
step 1: subtract 13x from both sides of the equation -3y=-13x+18
step 2: divide -3 on both sides of the equation y = 13/3x - 6
Therefore your slope intercept form of 13x - 3y = 18 is y = 13/3x - 6
Q2.Express the following equation in slope-intercept form: 17x - 10y = 80
Step 1: Subtract 17x from both sides of the equation -10y = -17x + 80
Step 2: Divide -10 from both sides of the equation y = 17/10x - 8
Therefore the slope intercept form of 17x - 10y = 80 is y =17/10x - 8
Q3.Express the following equation in slope-intercept form: 5x + 8y = 40
Step 1: subtract 5x from both sides of the equation 8y = -5x + 40
Step 2: divide 8 from both sides of the equation y = -5/8x + 5
Therefore the slope intercept form of 5x + 8y = 40 is y =-5/8x + 5